For all of the following use the five step process to solve the problem in the question, that is, state the null hypothesis, the alternative hypothesis, the level of significance and the type of test (z-test, t-test, p-value, confidence limit and whether it is two tailed, one-tailed upper or one tailed lower).
According to a recent survey, Americans get a mean of seven hours sleep per night. A random sample of 50 students a West Virginia University revealed the mean number of hours slept last night was 6 hours and 48 minutes (6.8 hours). The standard deviation of the sample was 0.9 hours. Is it reasonable to conclude that the students a West Virginia sleep less than the typical American?

Answer by stanbon(75887)

According to a recent survey, Americans get a mean of seven hours sleep per night.
A random sample of 50 students a West Virginia University revealed the mean number of hours slept last night was 6 hours and 48 minutes (6.8 hours).
-----
The standard deviation of the sample was 0.9 hours.
-------------
Is it reasonable to conclude that the students a West Virginia sleep less than the typical American?
Ho: u >= 7 hrs
Ha: u < 7 (claim)
--------------------------
z(6.8) = (6.8-7)/[0.9/sqrt(50)] = -1.5713
-----
p-value = P(z < -1.5713) = 0.0581
-----
Conclusion: Since the p-value is greater than 5%,
fail to reject Ho.
The test results to not support the claim.
==================
Cheers,
Stan H.
==================